# Evaluating Impacts of Traffic Regulations in Complex Mobility Systems Using Scenario-Based Simulations

### v1

@@ 5.1 Technical development
The implementation of the proposed system required work on two complementary fronts: the backend and the front end. On the backend side, the focus was on ensuring efficient computation and robust model execution, handling large-scale mobility and traffic data, and implementing the algorithms described in the previous sections. On the front end, the objective was to provide an accessible and intuitive interface that could be used by domain experts with varying levels of technical expertise. Together, these two components form an integrated technological system that balances computational rigor and usability, and enables both precise estimation of traffic outcomes and interactive policy experimentation.
@@
The front end was designed to ensure accessibility and usability for domain experts, including those without technical backgrounds. To this end, we developed an interactive dashboard that allows near real-time simulation of what-if scenarios under varying conditions. Users can directly modify key control parameters, described in Section 3 and 4, including policy fees per vehicle type, exempted vehicles, temporal windows of the policy application, behavioral parameters, and public transport settings. On the same page, simulation results are presented clearly and intuitively (see the next Section 5.3 for a wider discussion on qualitative and quantitative results). Global indicators summarize daily information for both the “as-is” baseline and “what-if” modified scenarios. They refer to indicators such as the total vehicle inflow per day, the maximum traffic within the day, total collected fees, and the estimated emission modification. For insight at a microscopic scale, interactive graphs are included to display the daily evolution of key quantities of interest in the current and modified scenarios. By changing the parameters, users can evaluate multiple scenarios corresponding to different combinations of parameters. The dashboard not only enables comparison between the baseline and modified scenarios, but also across different “what-if” scenarios (e.g., scenario 1 vs. scenario 2), supporting the assessment of alternative policy options.
@@ 5.2 Application: the case of Bologna, Italy
The proposed model was tested on a real-world case study in the city of Bologna in Italy, a major urban center with a large resident population and a complex multimodal mobility system. The city is already characterized by a limited traffic zone in the historical center. For experimental purposes, we simulated an additional circulation restriction applied within a broader zone: its northern boundary is delineated by the ring road and the A1-A14 motorway, while the southern boundary follows the city’s morphological constraints, namely the hilly area. Figure 3 represents the selected area. This configuration serves as an illustrative example of how the model can be applied to large-scale area-based policies.
@@
Bologna was further selected due to the advanced mobility data infrastructure maintained by the municipality. High-resolution data are collected from multiple sources, including traffic sensors, cameras, and vehicle registration data from the national motorization authority. These datasets are already used for real-time traffic monitoring, allowing us to estimate the current “as-is” traffic situation. In particular, the integration of multi-source data with statistical and data science preprocessing techniques enabled the estimation of inflows into the restricted area as well as trips originating within it, with a temporal discretization of $\nu=5$ minutes. Furthermore, we have estimates of the daily percentage of vehicles by Euro emission class currently circulating in the area. From Euro-0 to Euro-6 vehicles, respectively, the proportions are $0.059$, $0.012$, $0.034$, $0.054$, $0.198$, $0.176$, and $0.467$. Similarly, knowing the vehicular fleet composition in terms of segment, fuel type, and technology, we estimate the pollutant emissions per kilometer per vehicle typology in the area. We focus on the nitrogen oxides (NOx), a group of highly reactive gases emitted by vehicles that significantly contribute to air pollution. For each Euro-level class, the estimated emissions in grams per kilometer are: $0.2105847$, $0.217457$, $0.2401457$, $0.247239$, $0.135555$, $0.099559$, and $0.068246$.
@@ 5.3 Result assessment
In this section, we present the outcomes of the simulation scenarios under different access pricing policies and behavioural settings. To explore the potential effects of alternative parameter configurations, we construct a set of scenarios and systematically compare their impacts. The purpose of this analysis is not primarily to evaluate specific policies, but rather to illustrate a methodological approach for measuring variations in the model outputs under different conditions with the proposed methodology. The focus is on demonstrating how scenario-based simulations can be used to quantify and analyze changes in key performance indicators.
The comparison between scenarios is assessed by qualitative and quantitative analysis. First, we concentrate on a set of global metrics that summarize overall system behavior. They include the total number of vehicles entering in the area per day (i.e., daily inflow), the maximum number of vehicles circulating in the area during the day (i.e., max daily traffic), and the maximum number of vehicles circulating in the area during the policy activation hours (i.e., max traffic in policy hours). Note that these quantities differ in the measurement units: while inflows measure the cumulative number of entering vehicles during the whole day, the maximum traffic accounts for the maximum number of vehicles circulating at the same time in the interest period. Moreover, the analysis considers the total grams of NOx emitted by the vehicles circulating in the area during the day (i.e., daily emissions). Other than these key mobility indicators, other metrics are reported to better describe the what-if scenarios, such as the absolute number of vehicles changing their habits with a time-shift, with a modal-shift, or in another way (i.e., lost). Finally, the total earnings collected with the access tickets are summarized in the global indicator of daily revenue.
In addition to the baseline scenario, we consider a set of alternative scenarios designed to illustrate the effects of different pricing configurations and target populations. For example, we examine a daily fixed-price policy applied to the entire population during a 10h period (A1), in a longer 12h period (A2), and limited to morning hours (A3) and afternoon hours (A4). Moreover, we consider variations in which the price is differentiated by vehicle emission class with a focus on emissions reduction (A5) or differentiated per socio-economic level of the population to explore the effects of partial exemptions (A6). The parameters defining these scenarios are reported in Table 1. Furthermore, to assess the sensitivity of system outcomes to behavioral assumptions, we compare scenarios in which the same daily fixed-price policy is applied to 100% of the population while adjusting key behavioral parameters to capture extreme cases. These include a highly-reduced cost acceptability (B1), the removal of time-adjustment flexibility (B2), and the removal of modal shifts (B3). The parameters defining these scenarios are reported in Table 2.
@@
A3 Emission policy 8:00 AM 6:00 PM 0 {10 if $l=0,1,2,3$; 5 otherwise}
A4 Exempted policy 8:00 AM 6:00 PM 10 {5 $\forall l$}
A5 Afternoon policy 1:00 PM 6:00 PM 0 {5 $\forall l$}
A5 Morning policy 8:00 AM 1:00 PM 0 {5 $\forall l$}
Table 1: Parameter configuration of the scenarios B1–B4, assessing changes in policy impacts given changes in behavioral parameters. The remaining model parameters are set equal for all the scenarios: among behavioral parameters, $C_{50\%}=4-7$€, $\Delta_{ar50\%}=\Delta_{pr50\%}=1$ hours, $\Delta_{ar50\%}=\Delta_{pr50\%}=0.5$ hours, $\beta_{0}=-1.24$, $\beta_{1}=4.5$, $\beta_{2}=-1.45$, $\beta_{3}=-0.30$, $\beta_{4}=-0.034$; for the dwell time distribution, $\bar{\tau}=20$ minutes; for the emission conversion factor, $km\nu=2.5$
@@
B1 No economic-flexibility 0.1–0.5 1 1 True
@@
B3 No mode-flexibility 4–7 1 1 False
@@
Table 2: Parameter configuration of the scenarios B1–B3, assessing changes in policy impacts given changes in behavioral parameters. The remaining model parameters are set equal for all the scenarios: among control parameters, $t_{s}=$ 8:$0$ AM, $t_{e}=$ 6:$0$ PM, $C_{0}=5$€, $F_{e}=0\%$; among other behavioral parameters, $\Delta_{ar50\%}=1.5$ hours, $\Delta_{pr50\%}=1.5$ hours, $\beta_{0}=-1.24$, $\beta_{1}=4.5$, $\beta_{2}=-1.45$, $\beta_{3}=-0.30$, $\beta_{4}=-0.034$; for the dwell time distribution, $\bar{\tau}=20$ minutes; for the emission conversion factor, $km\nu=2.5$
@@
For the proposed scenarios A1–A6, Table 3 collects the measured value of these indicators. As expected, the results demonstrate that when access policies are activated, there is a significant reduction in inbound flows, traffic, and emissions. It is particularly interesting to assess how the intensity of these reductions varies depending on the applied policy. Scenarios A1–A4 highlight how the policy duration significantly affects traffic and emissions. Scenario A4 results in a larger reduction in daily inflow with respect to A1 (-56.5k vs -44.6k) and in traffic during policy hours (-5.0k vs -1.5k), with a more pronounced decrease in emissions (-209.8k NOx vs -166.7k NOx), but also a substantial reduction in maximum traffic during the day (-1.5k) in contrast to the increase in the A1 scenario (+1.4k). Scenarios A3 and A4 illustrate the effect of timing: the morning policy in A3 produces smaller impacts on inflow (-17.0k) and traffic (-0.6k policy-hours), whereas the afternoon policy in A4 modifies daily maximum traffic moderately (+0.9k) and maximum traffic in policy hours significantly (-4.4k), with emissions (-10.1% vs -11.9%) and revenue (112k in both scenarios) effects more balanced. Regarding policies tailored for Euro-class vehicles or including exemptions, daily inflow reductions range from -51.5k in A5 to -38.4k in A6, with A1 in between (-44.6k), while variations in traffic during policy hours are minor (-1.5k for A5, -1.4k for A6, -1.5k for A1). Emissions reductions follow the same pattern, from -192.8k NOx in A5 to -143.7k NOx in A6. Time- and mode-shifted vehicles, as well as lost vehicles, also show consistent trends across the three scenarios. Overall, the results indicate that despite slight numerical differences, the three scenarios produce broadly comparable impacts on system performance and revenue.
@@
Table 3: Comparison of global indicators in the modified scenarios of Table 1 with respect to the baseline situation
@@
Table 4: Comparison of global indicators in the modified scenarios of Table 2 with respect to the baseline situation
The qualitative analysis throughout visualizations is shown in Figures 4, 5, and 6. The higher daily maximum traffic in A1 compared to A4 is driven by the combined effect of vehicles anticipating or postponing trips, adding to an already high congestion level, whereas traffic remains more moderate throughout A4. Similarly, A1, A3, and A4 all experience peaks due to anticipatory and delayed trips, generating local surges in flow. Notably, morning traffic is lower overall, which explains why A3 shows smaller total reductions in daily inflow compared to A4. These patterns highlight how both the timing and duration of the policy shape daily traffic dynamics, congestion peaks, and the environmental effectiveness, beyond their impact on total demand. Figure 5 highlights the differences between scenarios A1, A5, and A6, which are present but largely homogeneous over time rather than exhibiting strong temporal patterns. This indicates that, in this case, global indicators are sufficient to capture and analyze the impacts of the scenarios, without requiring detailed time-resolved examination.
Figure 6 illustrates the distribution of the quantities over time for the scenarios B1–B3. The results reveal how behavioral constraints fundamentally reshape traffic patterns under mobility pricing. In scenario B1, traffic, inflow, and emissions become highly volatile with extreme peaks as travelers concentrate trips to avoid fees, followed by near-zero periods. Scenario B2, on the other hand, produces the most stable outcomes, with sustained moderate flows and emissions, and no high peaks near policy hours. The inability to shift from car to bus (scenario B3) generates an intermediate pattern with significant peaks close to policy activation hours, but maintained baseline traffic within policy hours. The examination of these scenarios highlights the strong dependence of system-level outcomes on behavioral flexibility assumptions, supporting the importance of comprehensive sensitivity analysis. In particular, small changes in travelers’ ability to reschedule trips or switch modes can lead to qualitatively different temporal traffic profiles and emission dynamics. These findings underscore that policy evaluations based on a single behavioral specification may be misleading and that robust mobility policy assessment requires explicit consideration of heterogeneous traveler responses.
Figure 4: Inflow [veh], traffic [veh], and emissions [g NOx] varying over time for the scenarios A1–A4 of Table 3
Figure 5: Inflow [veh], traffic [veh], and emissions [g NOx] varying over time for the scenarios A1, A5 and A6 of Table 3
Figure 6: Inflow [veh], traffic [veh], and emissions [g NOx] varying over time for the scenarios B1–B3 of Table 4

### v2

@@ 5.1 Technical development
The implementation of the proposed system requires work on two complementary fronts: the backend and the front end. On the backend side, the focus is on ensuring efficient computation and robust model execution, handling large-scale mobility and traffic data, and implementing the algorithms described in the previous sections. On the front end, the objective is to provide an accessible and intuitive interface that could be used by domain experts with varying levels of technical expertise. Together, these two components form an integrated technological system that balances computational rigor and usability, and enables both precise estimation of traffic outcomes and interactive policy experimentation.
@@
The frontend is designed to ensure accessibility and usability for domain experts, including those without technical backgrounds. To this end, we develop an interactive dashboard that allows near real-time simulation of what-if scenarios under varying conditions. Users can directly modify key control parameters, described in Section 3 and 4, including policy fees per vehicle type, exempted vehicles, temporal windows of the policy application, behavioral parameters, and public transport settings. On the same page, simulation results are presented clearly and intuitively (see the next Section 5.3 for a wider discussion on qualitative and quantitative results). Global indicators summarize daily information for both the “as-is” baseline and “what-if” modified scenarios. They refer to indicators such as the total vehicle inflow per day, the maximum traffic within the day, total collected fees, and the estimated emission modification. For insight at a microscopic scale, interactive graphs are included to display the daily evolution of key quantities of interest in the current and modified scenarios. By changing the parameters, users can evaluate multiple scenarios corresponding to different combinations of parameters. The dashboard not only enables comparison between the baseline and modified scenarios, but also across different “what-if” scenarios (e.g., scenario 1 vs. scenario 2), supporting the assessment of alternative policy options.
@@ 5.2 Application: the case of Bologna, Italy
The proposed model is tested on a real-world case study in the city of Bologna in Italy, a major urban center with a large resident population and a complex multimodal mobility system. The city is already characterized by a limited traffic zone in the historical center. For experimental purposes, we simulate an additional circulation restriction applied within a broader zone: its northern boundary is delineated by the ring road and the A1-A14 motorway, while the southern boundary follows the city’s morphological constraints, namely the hilly area. Figure 3 represents the selected area. This configuration serves as an illustrative example of how the model can be applied to large-scale area-based policies.
@@
Bologna is further selected due to the advanced mobility data infrastructure maintained by the municipality. High-resolution data are collected from multiple sources, including traffic sensors, cameras, and vehicle registration data from the national motorization authority. These datasets are already used for real-time traffic monitoring, allowing us to estimate the current “as-is” traffic situation. In particular, the integration of multi-source data with statistical and data science preprocessing techniques enables the estimation of inflows into the restricted area as well as trips originating within it, with a temporal discretization of $\nu=5$ minutes. Furthermore, we have estimates of the daily percentage of vehicles by Euro emission class currently circulating in the area. From Euro-0 to Euro-6 vehicles, respectively, the proportions are $0.059$, $0.012$, $0.034$, $0.054$, $0.198$, $0.176$, and $0.467$. Similarly, knowing the vehicular fleet composition in terms of segment, fuel type, and technology, we estimate the pollutant emissions per kilometer per vehicle typology in the area. We focus on the nitrogen oxides (NOx), a group of highly reactive gases emitted by vehicles that significantly contribute to air pollution. For each Euro-level class, the estimated emissions in grams per kilometer are: $0.2105847$, $0.217457$, $0.2401457$, $0.247239$, $0.135555$, $0.099559$, and $0.068246$.
@@ 5.3 Result assessment
In this section, we present the outcomes of the simulation scenarios under different access pricing policies and behavioural settings. To explore the potential effects of alternative parameter configurations, we construct a set of scenarios and systematically compared their impacts. The purpose of this analysis is not primarily to evaluate specific policies, but rather to illustrate a methodological approach for measuring variations in the model outputs under different conditions with the proposed methodology. The focus is on demonstrating how scenario-based simulations can be used to quantify and analyze changes in key performance indicators.
5.3.1 Scenario definition
In addition to the baseline scenario, we consider a set of alternative scenarios designed to illustrate the effects of different pricing configurations and target populations. These scenarios are selected to represent not only typical, realistic situations but also extreme cases that test the limit of the model. By including both ends of the spectrum, we can explore how the system responds under a variety of conditions and identify which parameters have the greatest influence on outcomes. This approach allows a systematic comparison across scenarios, providing insights into the sensitivity of the case study to the choice of parameter settings.
First, we propose a scenario analysis aimed at highlighting the effects of different policy timing schemes. We examine a daily fixed-price policy applied to the entire population during a 10h period from 8:00 AM to 6:00 PM (scenario A1), replicating standard timing schemes of a pricing policy against both morning and afternoon peak congestion. We then extend the policy hours to cover a longer period of 12h from 7:00 AM to 7:00 PM (scenario A2). Comparing the two addresses the sensitivity of the timing parameter in the study case by examining how shifting the time window by one hour affects system behavior (e.g., whether changes in inflow and traffic are equal, proportional, or exhibit a different pattern under such shifts). Next, we limit the policy to morning hours (8:00 AM to 1:00 PM, scenario A3) and to afternoon hours (1:00 PM to 6:00 PM, scenario A4). Comparing scenarios A3 and A4 allows us to understand whether the policy is more effective in alleviating traffic congestion during the morning peak or the afternoon peak.
Moreover, we incorporate scenarios where the policy targets different users based on their socio-economic characteristics, or given the differences in their vehicle emission levels. The goal of the comparison is to assess the ability of the policy to account for the differences among vehicle or user characteristics (e.g., socio-economic group, vehicle class) and to what extent the relevant indices are impacted by this group-differentiated policy. Focusing on emission reductions, we consider variations in which the price is differentiated by vehicle emission class, with a ticket fee of €10 for Euro-class levels 0–3 and a reduced fee of €5 for the remaining levels 4–6 (scenario A5). Exploring user exemptions, we differentiate per socio-economic level of the population and exempted users with economic fragility (i.e., yearly income <<13k euros), corresponding to 10% of the population (scenario A6).
To assess the sensitivity of system outcomes to behavioral assumptions, we compare scenarios in which the same daily fixed-price policy is applied to 100% of the users while adjusting key behavioral parameters to capture extreme cases. These include a highly-reduced cost acceptability, where the ticket fee is fixed at €5 but the acceptable price is reduced to be <<€0.5 (scenario B1). Then, we consider the removal of time-adjustment flexibility (scenario B2) and the removal of modal shifts (scenario B3), which prevent considering time- and mode- variability, respectively, as acceptable behavioral strategies. These three extreme cases each strongly disincentivize a specific behavioral strategy, thereby allowing us to isolate and analyze the effect of excluding that strategy. This approach is particularly informative because behavioral strategies are difficult to model and their associated parameters are often subject to substantial uncertainty. By examining the extreme case in which a given strategy is absent, we effectively explore a boundary state of the system.
The parameters defining the scenarios A1–A6 are reported in Table 1, while parameter configurations of scenarios B1–B3 are reported in Table 2.
@@ 5.3.1 Scenario definition
A3 Morning policy 8:00 AM 1:00 PM 0 {5 $\forall l$}
A4 Afternoon policy 1:00 PM 6:00 PM 0 {5 $\forall l$}
A5 Emission policy 8:00 AM 6:00 PM 0 {10 if $l=0,1,2,3$; 5 otherwise}
A6 Exempted policy 8:00 AM 6:00 PM 10 {5 $\forall l$}
Table 1: Parameter configuration of the scenarios A1–A6, assessing changes in policy impacts given changes in control parameters. The remaining model parameters are set equal for all the scenarios: among behavioral parameters, $C_{50\%}=4-7$€, $\Delta_{a50\%}=\Delta_{p50\%}=1$ hours, $\Delta_{ar50\%}=\Delta_{pr50\%}=1.5$ hours, $\beta_{0}=-1.24$, $\beta_{1}=4.5$, $\beta_{2}=-1.45$, $\beta_{3}=-0.30$, $\beta_{4}=-0.034$; for the dwell time distribution, $\bar{\tau}=20$ minutes; for the emission conversion factor, $km\nu=2.5$
@@
B1 No economic-flexibility 0.1–0.5 1.5 1.5 True
@@
B3 No mode-flexibility 4–7 1.5 1.5 False
@@
Table 2: Parameter configuration of the scenarios B1–B3, assessing changes in policy impacts given changes in behavioral parameters. The remaining model parameters are set equal for all the scenarios: among control parameters, $t_{s}=$ 8:$0$ AM, $t_{e}=$ 6:00 PM, $C_{0}=5$€, $F_{e}=0\%$; among other behavioral parameters, $\Delta_{ar50\%}=1.5$ hours, $\Delta_{pr50\%}=1.5$ hours, $\beta_{0}=-1.24$, $\beta_{1}=4.5$, $\beta_{2}=-1.45$, $\beta_{3}=-0.30$, $\beta_{4}=-0.034$; for the dwell time distribution, $\bar{\tau}=20$ minutes; for the emission conversion factor, $km\nu=2.5$
5.3.2 Evaluation setup
The comparison between scenarios is assessed by qualitative and quantitative analysis. First, we concentrate on a set of global metrics that summarize overall system behavior. They include the total number of vehicles entering the area per day (i.e., daily inflow), the maximum number of vehicles circulating in the area during the day (i.e., max daily traffic), and the maximum number of vehicles circulating in the area during the policy activation hours (i.e., max traffic in policy hours). Note that these quantities differ in the measurement units: while inflows measure the cumulative number of entering vehicles during the whole day, the maximum traffic accounts for the maximum number of vehicles circulating at the same time in the interest period. Moreover, the analysis considers the total grams of NOx emitted by the vehicles circulating in the area during the day (i.e., daily emissions). Other than these key mobility indicators, other metrics are reported to better describe the what-if scenarios, such as the absolute number of vehicles changing their habits with a time-shift, with a modal-shift, or in another way (i.e., lost). Finally, the total earnings collected with the access tickets are summarized in the global indicator of daily revenue.
5.3.3 Findings
@@
For the proposed scenarios A1–A6, Table 3 collects the measured value of the global indicators. As expected, the results demonstrate that when access policies are activated, there is a significant reduction in inbound flows, traffic, and emissions. It is particularly interesting to quantify how the intensity of these reductions varies depending on the applied policy. Scenarios A1–A4 highlight how the policy duration significantly affects traffic and emissions. Scenario A2 results in a larger reduction in daily inflow with respect to A1 (-56.5k vs -44.6k) and in traffic during policy hours (-5.0k vs -1.5k), with a more pronounced decrease in emissions (-209.8k NOx vs -166.7k NOx), but also a substantial reduction in maximum traffic during the day (-1.5k) in contrast to the increase in the A1 scenario (+1.4k). Scenarios A3 and A4 illustrate the effect of timing: the morning policy in A3 produces smaller impacts on inflow (-17.0k) and traffic (-0.6k policy-hours), whereas the afternoon policy in A4 modifies daily maximum traffic moderately (+0.9k) and maximum traffic in policy hours significantly (-4.4k), with emissions (-10.1% vs -11.9%) and revenue (112k in both scenarios) effects more balanced.
Regarding policies tailored for Euro-class vehicles or including exemptions, daily inflow reductions range from -51.5k in A5 to -38.4k in A6, with A1 in between (-44.6k), while variations in traffic during policy hours are minor (-1.5k for A5, -1.4k for A6, -1.5k for A1). Emissions reductions follow the same pattern, from -192.8k NOx in A5 to -143.7k NOx in A6. Time- and mode-shifted vehicles, as well as lost vehicles, also show consistent trends across the three scenarios. Overall, the results indicate that despite slight numerical differences, the three scenarios produce broadly comparable impacts on system performance and revenue.
@@
Table 3: Comparison of global indicators in the scenarios A1–A6 with respect to the baseline situation
@@
Table 4: Comparison of global indicators in the scenarios B1–B3 with respect to the baseline situation
The qualitative analysis throughout visualizations is shown in Figures 4, 5, and 6. The higher daily maximum traffic in A1 compared to A2 is driven by the combined effect of vehicles anticipating or postponing trips, adding to an already high congestion level, whereas traffic remains more moderate throughout A4. Similarly, A1, A3, and A4 all experience peaks due to anticipatory and delayed trips, generating local surges in flow. Notably, morning traffic is lower overall, which explains why A3 shows smaller total reductions in daily inflow compared to A4. These patterns highlight how both the timing and duration of the policy shape daily traffic dynamics, congestion peaks, and the environmental effectiveness, beyond their impact on total demand.
Figure 5 highlights the differences between scenarios A1, A5, and A6, which are present but largely homogeneous over time rather than exhibiting strong temporal patterns. This indicates that, in this case, global indicators are sufficient to capture and analyze the impacts of the scenarios, without requiring detailed time-resolved examination.
Figure 6 illustrates the distribution of the quantities over time for the scenarios B1–B3. The results reveal how behavioral constraints fundamentally reshape traffic patterns under mobility pricing. In scenario B1, traffic, inflow, and emissions become highly volatile with extreme peaks as travelers concentrate trips to avoid fees, followed by near-zero periods in policy activation hours. Scenario B2, on the other hand, produces the most stable outcomes, with sustained moderate flows and emissions, and no high peaks near policy hours. The inability to shift from car to bus (scenario B3) generates an intermediate pattern with significant peaks close to policy activation hours, but maintained baseline traffic within policy hours. The examination of these scenarios highlights the strong dependence of system-level outcomes on behavioral flexibility assumptions, supporting the importance of comprehensive sensitivity analysis. In particular, small changes in travelers’ ability to reschedule trips or switch modes can lead to qualitatively different temporal traffic profiles and emission dynamics. These findings underscore that policy evaluations based on a single behavioral specification may be misleading and that robust mobility policy assessment requires explicit consideration of heterogeneous traveler responses.
Figure 4: Inflow [veh], traffic [veh], and emissions [g NOx] varying over time for the scenarios A1–A4
Figure 5: Inflow [veh], traffic [veh], and emissions [g NOx] varying over time for the scenarios A1, A5–A6
Figure 6: Inflow [veh], traffic [veh], and emissions [g NOx] varying over time for the scenarios B1–B3